# FRIEDA: Benchmarking Multi-Step Cartographic Reasoning in Vision-Language Models

### v1

@@ E.2 Statistical Significance of FRIEDA-direct Results
As FRIEDA partitions questions into a large number of fine-grained categories, some subsets contain relatively few examples (fewer than 100). In such cases, raw accuracy comparisons can be unreliable due to limited sample size. To more rigorously assess whether observed performance differences within these smaller subcategories are statistically meaningful, we apply McNemar’s test c:mcnemar1947 on the top three proprietary models. We use the exact binomial version of the test when the number of disagreements is small (<50<50), and the $\chi$-squared version with correction when disagreements are larger ($\geq 50$). Table 7 reports the resulting $p$-values.
@@ F.2 Analyses on Model Size
In the main evaluation (Figure 3), the results deviate from the usual scaling law (c:kaplan2020scalinglaws), which states that the performance of the model improves with size. Among open-source models, LLaVA-NeXT, despite having the most parameters, ranks near the bottom, whereas Ovis-2.5-9B, the smallest model, ranks near the top. We, therefore, hypothesize that cartographic reasoning is not an emergent ability (i.e., a capability absent in smaller models but present in larger ones). To test this, we evaluate the InternVL3.5 family (c:wang2025internvl35) on FRIEDA: 1B, 2B, 4B, 8B, 14B, 30BA3B, 38B, 241BA28B where ‘A’ denotes parameters active at inference. The trend (Figure 19, Table 16) shows modest gains up to roughly 30B parameters, followed by degradation thereafter.
@@ G.1 Document and Infographics VQA
Benchmarks in this domain have established baselines for LVLM reasoning over structured text and graphical images, including charts. In the document domain, DocVQA (c:matthew2021docvqa) introduces a large-scale question-answering dataset over real forms and reports. DocVXQA (c:souibgui2025docvxqa) builds upon the benchmark to design a self-explanatory framework that produces interpretable rationales for LVLMs. Docopilot (c:duan2025docopilot) evaluates LVLMs on scientific articles, which not only test text understanding but also the interpretation of embedded figures such as charts. For graphics, InfographicsVQA (c:matthew2022infovqa) tests joint reasoning over text, layout, and pictorial elements in visually rich infographics. InfoChartQA (c:lin2025infochartqa) extends this by pairing plain charts and infographics and identifying design elements that degrade LVLM performance. In general, VQA evaluation on frontier LVLMs reveals a consistent trend: competence at high-level patterns, such as identifying trends in the data or the extrema, but struggles with precise value extraction and robustness. FRIEDA evaluates these shortcomings in a cartographic setting, where layout, symbols, legends, scales, and compass orientation interact tightly to measure how well LVLMs integrate these signals to answer map-based questions.
@@ G.2 Map VQA
MapQA (c:chang2022mapqa) establishes a baseline for choropleth map understanding by creating question-answer pairs targeting value retrieval and region identification. However, the dataset is limited to a single map type (i.e., choropleth maps) and geographically restricted to the United States, thereby limiting the diversity of cartographic styles and toponyms. MapWise (c:mukhopadhyay2025mapwise) broadens geographic coverage to the United States, India, and China and introduces diverse question templates for probing relative spatial relationships; yet, it still relies solely on choropleth maps and remains constrained to single-map reasoning, which limits its ability to model real-world cartographic complexity. MapIQ (c:srivastava2025mapiq) further advances visualization literacy by introducing cartograms and proportional-symbol maps, which are commonly used in analytical tasks. While the expanded map diversity is valuable, MapIQ’s maps are generated using map-coloring tools rather than drawn from heterogeneous, noisy real-world documents. In contrast, FRIEDA explicitly captures this real-world variability by sourcing map images directly from government and scientific reports.
@@
Benchmarks centered on navigation often require more complex reasoning than simple semantic retrieval, yet they tend to remain domain-narrow. MapEval (c:dihan2025mapeval) evaluates LVLMs’ geospatial reasoning through multiple-choice travel-planning questions spanning 180 cities. Still, it relies on standard web basemaps (e.g., Google Maps) whose clean, uniform designs lack the layered, domain-specific symbology (e.g., variable legends, irregular projections, and customized north arrows) often found in professional cartography. ReasonMap (c:feng2025canmllmguidemehome) moves beyond basemaps by using high-resolution transit maps and designing navigation tasks that closely simulate real-world subway routing, though its scope is restricted to transit systems. MapBench (c:xing2025mapbench) evaluates LVLMs’ spatial reasoning and chain-of-thought inference by testing outdoor navigation performance on diverse map types, such as park and trail maps. Despite their contributions, all of these benchmarks remain focused on navigation-centric tasks. In contrast, our benchmark generalizes spatial reasoning across six distinct spatial relations that extend well beyond navigation, capturing the broader landscape of cartographic reasoning required in professional and scientific contexts.
@@
Specialized benchmarks address domain-specific needs or particular visual modalities, but they tend to trade breadth for depth. PEACE (c:huang2025peace) introduces a benchmark focused on geologic map understanding and develops a framework for answering domain-specific questions, such as identifying lithologic units, fault lines, and structural patterns. While the benchmark and the approach is highly effective for geology-specific evaluation, the scope is limited to a single scientific domain, and it lacks the thematic diversity required for broader cartographic reasoning. CartoMark (c:zhou2024cartomark) provides a wide range of maps across various styles, but its core task centers on simple pattern recognition, such as scene classification and text annotation. These tasks primarily test perceptual recognition and, in many cases, do not require reasoning at all. ReMI (c:kazemi2025remi) offers a framework for multi-image reasoning that evaluates how models integrate and compare information across visual inputs. However, ReMI operates on natural images and uses simple web-based maps. Therefore, it does not assess the specialized challenges of multi-map cartographic reasoning, such as aligning heterogeneous legends, reconciling differing spatial scales, and interpreting mismatched orientations across maps. These capabilities form the core of FRIEDA’s multi-map setting, which reflects real-world analytical scenarios where experts must synthesize information from multiple, heterogeneous cartographic sources.
@@
MapQA (c:chang2022mapqa) 1 1
CartoMark (c:zhou2024cartomark) 13 7
MapWise (c:mukhopadhyay2025mapwise) 3 3
MapIQ (c:srivastava2025mapiq) 1 6
MapBench (c:xing2025mapbench) UNK 9
MapEval (c:dihan2025mapeval) 54 1
ReMi (c:kazemi2025remi) 100? 1
PEACE (c:huang2025peace) 2 1
ReasonMap (c:feng2025canmllmguidemehome) 13 1
@@
Note: ReMi (c:kazemi2025remi) reports counts by city, not by country; consequently, the corresponding country total is less than 100.
@@ G.3 Spatial Reasoning
Spatial reasoning benchmarks have advanced model capabilities in perception and localization, though often outside the cartographic domain. Benchmarks such as SpatialVLM (c:chen2024spatialvlm) and SpatialRGPT (c:cheng2024spatialrgpt) focus on natural images, testing a model’s ability to reason about 2D and 3D spatial relationships, relative positions, and object dimensions in photographic scenes. In the geospatial domain, GeoChain (c:yerramilli2025geochain) enhances tasks such as geolocation by inducing step-by-step geographic reasoning to link visual cues to geographic entities. However, these works primarily rely on natural scene understanding or semantic knowledge retrieval and do not engage with the abstract symbolic conventions unique to maps. FRIEDA closes this gap by evaluating multi-step cartographic reasoning, in which models must not only perceive space but also decode specific symbolic rules to infer topological, metric, and directional relations across heterogeneous real-world maps.

### v2

@@ E.2 Statistical Significance of FRIEDA-direct Results
As FRIEDA partitions questions into a large number of fine-grained categories, some subsets contain relatively few examples (fewer than 100). In such cases, raw accuracy comparisons can be unreliable due to limited sample size. To more rigorously assess whether observed performance differences within these smaller subcategories are statistically meaningful, we apply McNemar’s test McNEMAR (1947) on the top three proprietary models. We use the exact binomial version of the test when the number of disagreements is small (<50<50), and the $\chi$-squared version with correction when disagreements are larger ($\geq 50$). Table 7 reports the resulting $p$-values.
@@ F.2 Analyses on Model Size
In the main evaluation (Figure 3), the results deviate from the usual scaling law (Kaplan et al., 2020), which states that the performance of the model improves with size. Among open-source models, LLaVA-NeXT, despite having the most parameters, ranks near the bottom, whereas Ovis-2.5-9B, the smallest model, ranks near the top. We, therefore, hypothesize that cartographic reasoning is not an emergent ability (i.e., a capability absent in smaller models but present in larger ones). To test this, we evaluate the InternVL3.5 family (Wang et al., 2025b) on FRIEDA: 1B, 2B, 4B, 8B, 14B, 30BA3B, 38B, 241BA28B where ‘A’ denotes parameters active at inference. The trend (Figure 19, Table 16) shows modest gains up to roughly 30B parameters, followed by degradation thereafter.
@@ G.1 Document and Infographics VQA
Benchmarks in this domain have established baselines for LVLM reasoning over structured text and graphical images, including charts. In the document domain, DocVQA (Mathew et al., 2021) introduces a large-scale question-answering dataset over real forms and reports. DocVXQA (Souibgui et al., 2025) builds upon the benchmark to design a self-explanatory framework that produces interpretable rationales for LVLMs. Docopilot (Duan et al., 2025) evaluates LVLMs on scientific articles, which not only test text understanding but also the interpretation of embedded figures such as charts. For graphics, InfographicsVQA (Mathew et al., 2022) tests joint reasoning over text, layout, and pictorial elements in visually rich infographics. InfoChartQA (Lin et al., 2025) extends this by pairing plain charts and infographics and identifying design elements that degrade LVLM performance. In general, VQA evaluation on frontier LVLMs reveals a consistent trend: competence at high-level patterns, such as identifying trends in the data or the extrema, but struggles with precise value extraction and robustness. FRIEDA evaluates these shortcomings in a cartographic setting, where layout, symbols, legends, scales, and compass orientation interact tightly to measure how well LVLMs integrate these signals to answer map-based questions.
@@ G.2 Map VQA
MapQA (Chang et al., 2022) establishes a baseline for choropleth map understanding by creating question-answer pairs targeting value retrieval and region identification. However, the dataset is limited to a single map type (i.e., choropleth maps) and geographically restricted to the United States, thereby limiting the diversity of cartographic styles and toponyms. MapWise (Mukhopadhyay et al., 2025) broadens geographic coverage to the United States, India, and China and introduces diverse question templates for probing relative spatial relationships; yet, it still relies solely on choropleth maps and remains constrained to single-map reasoning, which limits its ability to model real-world cartographic complexity. MapIQ (Srivastava et al., 2025) further advances visualization literacy by introducing cartograms and proportional-symbol maps, which are commonly used in analytical tasks. While the expanded map diversity is valuable, MapIQ’s maps are generated using map-coloring tools rather than drawn from heterogeneous, noisy real-world documents. In contrast, FRIEDA explicitly captures this real-world variability by sourcing map images directly from government and scientific reports.
@@
Benchmarks centered on navigation often require more complex reasoning than simple semantic retrieval, yet they tend to remain domain-narrow. MapEval (Dihan et al., 2025) evaluates LVLMs’ geospatial reasoning through multiple-choice travel-planning questions spanning 180 cities. Still, it relies on standard web basemaps (e.g., Google Maps) whose clean, uniform designs lack the layered, domain-specific symbology (e.g., variable legends, irregular projections, and customized north arrows) often found in professional cartography. ReasonMap (Feng et al., 2025) moves beyond basemaps by using high-resolution transit maps and designing navigation tasks that closely simulate real-world subway routing, though its scope is restricted to transit systems. MapBench (Xing et al., 2025) evaluates LVLMs’ spatial reasoning and chain-of-thought inference by testing outdoor navigation performance on diverse map types, such as park and trail maps. Despite their contributions, all of these benchmarks remain focused on navigation-centric tasks. In contrast, our benchmark generalizes spatial reasoning across six distinct spatial relations that extend well beyond navigation, capturing the broader landscape of cartographic reasoning required in professional and scientific contexts.
@@
Specialized benchmarks address domain-specific needs or particular visual modalities, but they tend to trade breadth for depth. PEACE (Huang et al., 2025) introduces a benchmark focused on geologic map understanding and develops a framework for answering domain-specific questions, such as identifying lithologic units, fault lines, and structural patterns. While the benchmark and the approach is highly effective for geology-specific evaluation, the scope is limited to a single scientific domain, and it lacks the thematic diversity required for broader cartographic reasoning. CartoMark (Zhou et al., 2024) provides a wide range of maps across various styles, but its core task centers on simple pattern recognition, such as scene classification and text annotation. These tasks primarily test perceptual recognition and, in many cases, do not require reasoning at all. ReMI (Kazemi et al., 2025) offers a framework for multi-image reasoning that evaluates how models integrate and compare information across visual inputs. However, ReMI operates on natural images and uses simple web-based maps. Therefore, it does not assess the specialized challenges of multi-map cartographic reasoning, such as aligning heterogeneous legends, reconciling differing spatial scales, and interpreting mismatched orientations across maps. These capabilities form the core of FRIEDA’s multi-map setting, which reflects real-world analytical scenarios where experts must synthesize information from multiple, heterogeneous cartographic sources.
@@
MapQA (Chang et al., 2022) 1 1
CartoMark (Zhou et al., 2024) 13 7
MapWise (Mukhopadhyay et al., 2025) 3 3
MapIQ (Srivastava et al., 2025) 1 6
MapBench (Xing et al., 2025) UNK 9
MapEval (Dihan et al., 2025) 54 1
ReMi (Kazemi et al., 2025) 100? 1
PEACE (Huang et al., 2025) 2 1
ReasonMap (Feng et al., 2025) 13 1
@@
Note: ReMi (Kazemi et al., 2025) reports counts by city, not by country; consequently, the corresponding country total is less than 100.
@@ G.3 Spatial Reasoning
Spatial reasoning benchmarks have advanced model capabilities in perception and localization, though often outside the cartographic domain. Benchmarks such as SpatialVLM (Chen et al., 2024a) and SpatialRGPT (Cheng et al., 2024) focus on natural images, testing a model’s ability to reason about 2D and 3D spatial relationships, relative positions, and object dimensions in photographic scenes. In the geospatial domain, GeoChain (Yerramilli et al., 2025) enhances tasks such as geolocation by inducing step-by-step geographic reasoning to link visual cues to geographic entities. However, these works primarily rely on natural scene understanding or semantic knowledge retrieval and do not engage with the abstract symbolic conventions unique to maps. FRIEDA closes this gap by evaluating multi-step cartographic reasoning, in which models must not only perceive space but also decode specific symbolic rules to infer topological, metric, and directional relations across heterogeneous real-world maps.